OpenStudy (taytay):

How do you find the inflection point of f(s) = 2s^3+18s^2+5

OpenStudy (anonymous):

differentiate f(s) wif respect to s. at inflection point, second derivative equal to zero.

OpenStudy (taytay):

so I have (-6, 0) is that right?

OpenStudy (anonymous):

i got s=-3

OpenStudy (anonymous):

(-3,-97)

OpenStudy (amistre64):

y' = 6s^2 + 36s y'' = 12s y'' = 0 when x = 0 if I did it right

OpenStudy (amistre64):

and I didnt

OpenStudy (amistre64):

y'' = 12s +36 :) s= -36/12 = -3

OpenStudy (amistre64):

f(-3) = 2(-3)^3+18(-3)^2+5 f(-3) = 2(-27) +18(9) +5 f(-3) = -54 +162 +5 f(-3) = -54 +167 = 113 I hope :)

OpenStudy (anonymous):

the cord should be (-3,113)

OpenStudy (anonymous):

you are right